if c is positive and 2ax2+3bx+5c=0 does not have any real roots,then prove that 2a-3b+5c>0.


Thanks!

Grade:11

2 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

 

Dear Palak J

Ans: discriminant part is >0 and hence (3b)^2 < 4*a*c and hence 3b>2(10a*c)^(0.5)

now 2a-3b+5c > 2a-2(10a*c)^(0.5)+5c

or 2a-3b+5c>(p+q)^2

where p=(2a)^(0.5)   & q=(5c)^(0.5)

As this is a square term and hence is>0

hence 2a-3b+5c>0(proved)

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Palak J !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur

mycroft holmes
14 years ago

Since the given quadratic has no real roots, either f(x)>0 for all x or f(x)<0 for all x.

 

We are given that 5c = f(0)>0

 

Hence 2a-3b+5c = f(-1)>0